the sum of present ages of a father and son is 53 years. four years ago, the father's age was 4 times the of the son. find their present ages.
Answer by TimothyLamb(4379)

x = father age now
y = son age now
---
x + y = 53
x - 4 = 4(y - 4)
---
x + y = 53
x - 4 = 4y - 16
---
x + y = 53
x - 4y = -12
---
put the system of linear equations into standard form
---
x + y = 53
x - 4y = -12
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = father age now = 40
y = son age now = 13
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php